OpenStudy (anonymous):

write the expression in simplest radical form: -2x√72x^3y^4

OpenStudy (anonymous):

\[-2x\sqrt{72x^3y^4}\] Is it like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it might be easier to split up the radical like this: \[\sqrt{72}\sqrt{x^3}\sqrt{y^4}\] which are much easier to simplify.